OpenStudy (anonymous):

There are six black socks, six blue socks, 12 white socks, and four patterned socks in the drawer. What is the probability of pulling a blue sock, replacing it, and then drawing a patterned sock out of the drawer? Last question for the day my brain is done. I know this isn't hard but I just can't get it.

OpenStudy (anonymous):

probability is just how many divided by total number. For blue, it's 6 (blue ones) divided by (6 black+6 blue +12 white + 4 patterned) For patterned is 4(patterned) divided by total. BUT since you want the probability of BOTH things happening you just multiply the two results.

OpenStudy (anonymous):

P(blue socks) = 6/28 = 3/14 since socks is replaced, the total remains 28 so P(patterned sock) = 4/28 = 1/7

OpenStudy (anonymous):

6/28 TIMES 4/28 or 24/784 which I'm sure you can simplify more.

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

6/28 TIMES 4/28 or 24/784

OpenStudy (anonymous):

3/10 13/60 2/45 4/9 POSSIBLE ANSWERS